OpenStudy (anonymous):

Find the limit of (1+sec(x))/tan(x) as x approaches to pi/2.

OpenStudy (tkhunny):

Do youget l'Hopital?

OpenStudy (raffle_snaffle):

tan=sin/cos?

OpenStudy (anonymous):

Yes. But do I keep doing derivatives over and over again?

OpenStudy (raffle_snaffle):

you need to reduce the sec/tan

OpenStudy (raffle_snaffle):

I am not sure if reduce is the right word...

OpenStudy (tkhunny):

1) Do you KNOW it is an Indeterminate Form? If so, what type is it?

OpenStudy (anonymous):

Do I try to find the common denominator for the numerator?

OpenStudy (raffle_snaffle):

are you using 'Hopital rule/

OpenStudy (raffle_snaffle):

if so you need to detmerine if it's in indeterminate form

OpenStudy (tkhunny):

Ni, the '1' is not significant if it is of type \(\infty/\infty\).

OpenStudy (raffle_snaffle):

or 0/0

OpenStudy (tkhunny):

The 1 would be significant if it were 0/0 without it. That would make it 1/0 and NOT an indeterminate form.

OpenStudy (anonymous):

So I got (cosx+1)/sin(x)=-sinx/cosx=-cosx/-sinx=cosx/sinx,

OpenStudy (anonymous):

I get it now.

OpenStudy (tkhunny):

Don't do that. At pi/2, this is an indeterminate form of type \(\infty/\infty\). \(\dfrac{\dfrac{d}{dx}(1 + \sec(x))}{\dfrac{d}{dx}\tan(x)} = \dfrac{\sec(x)\tan(x)}{\sec^{2}(x)} = \dfrac{\tan(x)}{\sec(x)}\) You need to learn how to work with tangent and secant directly. Now what?

OpenStudy (raffle_snaffle):

tkhunny what is your anser?

OpenStudy (tkhunny):

We are waiting for Idealist to pipe in.

OpenStudy (raffle_snaffle):

you can solve it without using the "rule"

OpenStudy (raffle_snaffle):